Discuss the analytical foundations of the Phillips curve. Explain how an unemployment rate of 5% may be inflationary in one country, whilst deflationary in another
Answer #1

The Phillips curve shows the relationship between unemployment and inflation. It was originally started with a relationship between wages and unemployment. The curve was adapted as a change in wages meant an increase in labor cost that increased total costs for the employer.

Pence level SRAS До, — , У, У, pop и, ч, и,

From the graph we can see that there is a trade off between unemployment and inflation.If, for example, consumption increases, this will lead to aggregate demand to shift right from AD1 to AD2. This will increase the price level from p1 to p2 and also increase GDP. When GDP increase more factors of productions are being used, so unemployment falls from u1 to u2, and with an increase in price, an increase in the inflation rate will also occur, show by the move from i1 to i2. The same works if consumption falls, aggregate demand will also fall, shifting the curve left from AD1 to AD3. With this shift price falls to p3 and output fall to Y3. As output fall less factors of production are being utilised so unemployment rises to u3.

The original formula for the Phillips curve above was -

πt=πte+(μ+z)−α ut

But, Inflation before 1970 was almost zero so πte that is expected inflation was considered to be 0. So, the equation can be written as -

πt=(μ+z)−α πt

The Phillips Curve Equation can also be derived by aggregate supply relation -

P = Pe(1+μ)(1-αu+z)

Firstly, add time subscripts to the price level and the expected price level then divide whole equation by last year's price level

Pt / Pt-1 = P^et / Pt-1 (1+μ)(1-αu+z)

Rewrite the left hand-side of the equation as -

Pt / Pt-1 = Pt - Pt-1 + Pt-1 / Pt-1 = 1 + Pt - Pt-1 / Pt-1 = 1 + πt

Then, doing the same for the expected price level and substituting the new expressions into the equation divide through by one plus the expected inflation rate and one plus the mark-up

(1 + πt) / (1 + π^et)(1+μ) = 1-αu+z

Finally, the left-hand side of the equation can be approximated providing the values are not too large and re-arranged to give the equation for the Phillips Curve

πt=πte+(μ+z)−αut